OpenStudy (anonymous):

Find the general solution to the ODE:\[y''+9y=3t+e^{3t}\]I have the answer but I don't know how to do this one

OpenStudy (anonymous):

the answer i have in my exam review is \[y=\frac{t}{3}+\frac{e^{3t}}{18}+c_{1}sin(3t)+c_{2}cos(3t)\]and I'm just wondering how to do it because my professor only gave us the answer and I don't know how to start

OpenStudy (accessdenied):

Do you know how to solve the equation for a single function: Like solving each of these: \( y" + 9y = 3t \), \(y"+ 9y = e^{3t} \)

OpenStudy (anonymous):

ok so all I have to do is solve it separately and then combine them in the end?

OpenStudy (accessdenied):

Essentially, yes. There is something called the Principle of Superposition that states that, to find the particular solution of \(y" + p(x) y' + q(x) y = f_1 (x) + f_2 (x) + ... + f_N (x) \) If \(Y_j\) is a solution to \(y" + p(x) y'+ q(x) y = f_J(x) \), then \(Y_1 + Y_2 + Y_3 + ... + Y_N\) is a particular solution to the entire equation.

OpenStudy (anonymous):

ok this makes a lot of sense now. thank you so much. i know how to do the first part, but i am not too familiar with the \(e^{3t}\) portion. can you guide me as to how to start?

OpenStudy (accessdenied):

Hmm.. admittedly, I am rusty on these second order DE. So, I gotta use a book. :P Well, I believe there are a few methods that we could use to find the particular solution here. Do you use Variation of Parameters or Undetermined Coefficients typically? \( y"+ 9y = e^{3t} \)

OpenStudy (accessdenied):

And I think undetermined coefficients would be easier here. If we consider that the derivative of e^(3t) is always some Ae^(3t), we could try y=Ae^(3t) and solve for A...

OpenStudy (anonymous):

OK if I do that then I would end up with \(y=\frac{e^{3t}}{18}\) and combining this and the \(y=\frac{t}{3}\) from the first part I still don't know how to get the cos and sin part in the last half of the answer. Do I solve \(y''+9y=0\) in order to get them?

OpenStudy (accessdenied):

We get the particular solution out of these methods, so the general solution is just tacking on the Homogenous DE's solution. Yes. :)

OpenStudy (anonymous):

so I need to solve 3 ODE's to get this answer, in short? \[y''+9y=3t\]\[y''+9y=e^{3t}\]and \[y''+9y=0\]and then I will get the answer I want? Thanks and sorry for keeping you so long.

OpenStudy (accessdenied):

Yep. If we think about it, by combining all those together: \( y = Y_{p1} + Y_{p2} + y_H \) y_(p1) is solution to 1 DE, y_(p2) to another, and Y_H the homogenous DE solution The original DE splits up : \(y" + p(x) y'+ q(x) y = f(x) \\ \to (Y_{p1} + Y_{p2} + y_H)" + p(x) (Y_{p1} + Y_{p2} + y_H)' + q(x) (Y_{p1} + Y_{p2} + y_H) = f(x) \\ \to (y_H" + p(x)y_H' + q(x) y_H) + (Y_{p1}" + p(x) Y_{p1}' + q(x) Y_{p1}) \\ \qquad + (Y_{p2}" + q(x) Y_{p2}' + p(x)Y_{p2}) \) The homogenous case will eliminate to 0, and those other DE we solved earlier will sum up to the other side's f(x) Sorry this took a while to write. I thought it was a strong point to understand though.

OpenStudy (anonymous):

that's ok and thank you so much this helps a lot. Let me attempt some more problems.

OpenStudy (accessdenied):

You're welcome! :)